OpenStudy (anonymous):

Peter invested $9,000 in a program that compounded monthly and after 6 years he had $15,000. What was the interest rate?

OpenStudy (anonymous):

do u know the formula of compound interest

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im doing something wrong tho i dont know whAT

OpenStudy (anonymous):

ok ok wait

OpenStudy (anonymous):

?

OpenStudy (anonymous):

see here i would directly solve this by putting the values given

OpenStudy (anonymous):

ok could you work this out plz so i can see

OpenStudy (anonymous):

given that time is 6 years and principal is $9,000

OpenStudy (anonymous):

and also Amount is given $15,000

OpenStudy (anonymous):

15000=9000(1+r/12)^72

OpenStudy (anonymous):

yes yes right

OpenStudy (anonymous):

6000=(1+r/12)^72

OpenStudy (anonymous):

now after this what?

OpenStudy (anonymous):

wait u r going wrong it should not be r/12 it should be r/1200 and then n would be 6

OpenStudy (anonymous):

6000=(1+r/1200)^6??

OpenStudy (anonymous):

yes now do it

OpenStudy (anonymous):

how u get 1200

OpenStudy (anonymous):

way to big of number that cant be right